# A refined Weyl character formula for comodules on $\operatorname{GL}(2,A)$

### v2

@@ 8. Duals of symmetric powers and symmetric tensors
Let $k[G]:=A[x_{ij},1/D]$ with $G:=\operatorname{Spec}(k[G])\cong\operatorname{GL}(2,A)$. Let $\Delta:W\rightarrow W\otimes_{A}k[G]$ be a comodule on $G$. Let $\Delta_{G}$ be the comultiplication on $G$. For any commutative $A$-algebra $B$ there is for any $g\in h_{G}(B)$ a commutative diagram
@@
is a direct sum of $k[N]$-comodules. It is the refined weight space decomposition of $\operatorname{Sym}^{2}(V)^{*}$. We also see directly that $\operatorname{Sym}^{2}(V)^{*}\cong\operatorname{S}_{2}(V)^{*}$ are isomorphic as $k[N]$-comodules by inpecting the structure matrices - they are ”equal”. If $A:=\mathbf{Z}$ there is no isomorphism $\operatorname{Sym}^{2}(V)^{*}\cong\operatorname{S}_{2}(V)^{*}$ of $k[\operatorname{GL}(2,\mathbf{Z})]$-comodules since this implies there is (taking double dual) an isomorphism $\operatorname{Sym}^{2}(V)\cong\operatorname{S}_{2}(V)$ which contradicts Theorem 7.3.
@@ 9. The adjoint representation in coordinates
In this section we give a direct and elementary construction of the adjoint representation of $\operatorname{GL}(2,A)$ using the coordinate ring and the algebra of distributions. We recover the explicit formula for the adjoint representation and the refined weight space decompostion constructed in Section 2.
@@
Let $z_{11},z_{12}\,z_{21},z_{22}$ be independent coordinates over $A$ and let $A[z_{ij}]$ be the polynomial ring in the $z_{ij}$. Let $G:=\operatorname{GL}(2,A)$. Define the following actions:
@@
$\phi:(I/I^{2})^{*}\cong\operatorname{Dist}^{+}_{1}(\operatorname{GL}(2,A),e):=\operatorname{Lie}(\operatorname{GL}(2,A))\cong\mathfrak{gl}(2,A)$
@@
$\operatorname{Ad}^{*}:\operatorname{Lie}(\operatorname{GL}(2,A))\rightarrow\operatorname{Lie}(\operatorname{GL}(2,A))\otimes k[G]$
@@
$\operatorname{Ad}_{N}^{*}:\operatorname{Lie}(\operatorname{GL}(2,A))\rightarrow\operatorname{Lie}(\operatorname{GL}(2,A))\otimes k[N]$
@@
where the last matrix is the one from Section 2 where we defined the adjoint representation using the functor of points. Hence the map $\phi:(I/I^{2})^{*}\rightarrow\mathfrak{gl}(2,A)$ defined by $\phi(T_{1}):=e_{11},\phi(x):=e_{12},\phi(y):=e_{21},\phi(T_{2}):=e_{22}$ is an isomorphism of comodules.
When we pass to the diagonal torus we get a formula for the weights of $\operatorname{Lie}(\operatorname{GL}(2,A))$:
@@
Hence the vectors $T_{1},x,y,T_{2}$ have weights $\{(0,0),(-1,1),(1,-1),(0,0)\}$. We also recover the roots of $\operatorname{Lie}(\operatorname{GL}(2,A))$ as $R_{G}:=\{(-1,1),(1,-1)\}$.
@@
You may also compare the above calculation with the calculation of the adjoint representation of $\operatorname{SL}(2,A)$ on $\mathfrak{sl}(2,A)$ from Section 2.

### v4

@@ 9. Duals of symmetric powers and symmetric tensors
Let $k[G]:=A[x_{ij},1/D]$ with $G:=\operatorname{Spec}(k[G])\cong\operatorname{GL}_{2,A}$. Let $\Delta:W\rightarrow W\otimes_{A}k[G]$ be a comodule on $G$. Let $\Delta_{G}$ be the comultiplication on $G$. For any commutative $A$-algebra $B$ there is for any $g\in h_{G}(B)$ a commutative diagram
@@
is a direct sum of $k[N]$-comodules. It is the refined weight space decomposition of $\operatorname{Sym}^{2}(V)^{*}$. We also see directly that $\operatorname{Sym}^{2}(V)^{*}\cong\operatorname{S}_{2}(V)^{*}$ are isomorphic as $k[N]$-comodules by inpecting the structure matrices - they are ”equal”. If $A:=\mathbf{Z}$ there is no isomorphism $\operatorname{Sym}^{2}(V)^{*}\cong\operatorname{S}_{2}(V)^{*}$ of $k[\operatorname{GL}_{2,\mathbf{Z}}]$-comodules since this implies there is (taking double dual) an isomorphism $\operatorname{Sym}^{2}(V)\cong\operatorname{S}_{2}(V)$ which contradicts Theorem 8.3.
@@ 10. The adjoint representation in coordinates
In this section we give a direct and elementary construction of the adjoint representation of $\operatorname{GL}_{2,A}$ using the coordinate ring and the algebra of distributions. We recover the explicit formula for the adjoint representation and the refined weight space decompostion constructed in Section 2.
@@
Let $z_{11},z_{12}\,z_{21},z_{22}$ be independent coordinates over $A$ and let $A[z_{ij}]$ be the polynomial ring in the $z_{ij}$. Let $G:=\operatorname{GL}_{2,A}$. Define the following actions:
@@
$\phi:(I/I^{2})^{*}\cong\operatorname{Dist}^{+}_{1}(\operatorname{GL}_{2,A},e):=\operatorname{Lie}(\operatorname{GL}_{2,A})\cong\mathfrak{gl}_{2,A}$
@@
$\operatorname{Ad}^{*}:\operatorname{Lie}(\operatorname{GL}_{2,A})\rightarrow\operatorname{Lie}(\operatorname{GL}_{2,A})\otimes k[G]$
@@
$\operatorname{Ad}_{N}^{*}:\operatorname{Lie}(\operatorname{GL}_{2,A})\rightarrow\operatorname{Lie}(\operatorname{GL}_{2,A})\otimes k[N]$
@@
where the last matrix is the one from Section 2 where we defined the adjoint representation using the functor of points. Hence the map $\phi:(I/I^{2})^{*}\rightarrow\mathfrak{gl}_{2,A}$ defined by $\phi(T_{1}):=e_{11},\phi(x):=e_{12},\phi(y):=e_{21},\phi(T_{2}):=e_{22}$ is an isomorphism of comodules.
When we pass to the diagonal torus we get a formula for the weights of $\operatorname{Lie}(\operatorname{GL}_{2,A})$:
@@
Hence the vectors $T_{1},x,y,T_{2}$ have weights $\{(0,0),(-1,1),(1,-1),(0,0)\}$. We also recover the roots of $\operatorname{Lie}(\operatorname{GL}_{2,A})$ as $R_{G}:=\{(-1,1),(1,-1)\}$.
@@
You may also compare the above calculation with the calculation of the adjoint representation of $\operatorname{SL}_{2,A}$ on $\mathfrak{sl}_{2,A}$ from Section 2.